Anna consumes two goods: cricket games and fries. Her income is $60, the price of cricket games is
$20, and the price of a serving of fries is $10. Consider this table of marginal utilities:
Quantity MU from cricket
games

MU from fries

1 200 100
2 150 75
3 100 50
4 50 20
5 25 10
6 20 5

a) Complete the table above. [2 marks]

b) What is the Anna’s opportunity cost of going to a cricket game? [2 marks]

c) Suppose that you consumed 1 football game and 4 servings of fries. Why is this not the
equilibrium? [3 marks]

d) What is the equilibrium?
1
Expert's answer
2020-09-06T17:28:48-0400

a)


b) $10

c)

"20+4\times 10=60"

but


"\frac {MU_x}{p_x}=10"


"\frac {MU_y}{p_y}=2"

d)


"20x+10y=60"

"\frac {MU_x}{p_x}=\frac {MU_y}{p_y}"

x=2, y=2